OpenStudy (anonymous):

The sum of two numbers is 2014, while their difference is 42. what are the two number?

Directrix (directrix):

Let x and y be the two numbers. sum of two numbers is 2014 1) x + y = 2014 their difference is 42 2) x - y = 42

Directrix (directrix):

1) x + y = 2014 2) x - y = 42 -------------- Add the two equations 2x = 2056 Solve for x x = ? Then substitute the value of x into x + y = 2014 and solve for y.

OpenStudy (anonymous):

I typically do these types of problems in two parts. First, solve for one of the variables. i.e. x = y + 42. Then plug that into the other equation for x. So, x+y=2013 becomes y+42+y=2014. From there you can solve for both variables.